Suppose the interest rate is 9.8 % APR with monthly compounding. What is the present value of an annuity that pays $ 80 every three months for five ​years? ​(Note: Be careful not to round any intermediate steps less than six decimal​ places.)

The present value of the annuity is? (Round to the nearest​ cent.
Answer #1

APR = 9.8%

3-Month compounding rate = 9.8%*(3/12) = 2.45%

Number of compoundings = (5*12)/3 = 60/3 = 20

Payment every 3 months = $80

I = 2.45%; N = 20; PMT = 80, solve for PV.

PV = 1,253.04